OpenStudy (anonymous):

in a scientific notation (62*10^4)*(3.3*x10^2) equals i have no idea

OpenStudy (aaronq):

62*10^4=620000 3.3*x10^2=330 (62*10^4)*(3.3*x10^2)=620000*330

OpenStudy (ookawaiioo):

6.2*10^5 x 3.3*10^2 = (6.2*3.3)*10^7 =